# Quantum Signatures of Chaos from Free Probability

### v1

@@ 7.1 Future directions
  • In this work, we have examined the free convolution prediction for the sum of two free operators, focusing on spin-1/2 and spin-1 operators. While these results, specially, the spin-1/2 case, apply to a broad class of quantum mechanical systems, including spin chains and SYK-like models, it is crucial to extend the free convolution prediction to sums of other types of operators that naturally arise in quantum mechanics. In Sec. 3.2.1, we numerically determine the free additive convolution for the sum of two free spin-3/2, spin-2 and spin-5/2 operators as a first step in this direction. However, many important cases remain unexplored, such as bosonic operators in, say, the Bose-Hubbard model, and the position $x$ and momentum $p$ operators in quantum billiard systems, and numerous others. A systematic classification of free convolution predictions for different classes of operators would be highly valuable for future studies on quantum chaos and asymptotic freeness in quantum mechanics.
@@
  • In this paper we have considered spectral properties of the sum of two non-commutative operators (say $A$ and $B$). However, one can similarly consider the spectral properties of the products of two such operators, as well as their different linear combinations. If these variables are traceless, i.e., $\varphi(A)=\varphi(B)=0$, then their free product is trivial since the trace of a combination of them vanishes from the definition freeness, $\varphi\big{(}(AB)^{k}\big{)}=\varphi\big{(}ABAB\cdots AB\big{)}=0$. For variables with non-zero trace, one can obtain the free multiplicative convolution of these variables by using an analogue of the $R$-transform, known as the $S$-transform, which has the property that for a pair of free variables, it is multiplicative voiculescu1987multiplication . This allows us to compute the free multiplicative convolution of the eigenvalue density of these two operators and obtain the eigenvalue density of the product by using the rule $S_{\rho_{A}\boxtimes\rho_{B}}(x)=S_{\rho_{A}}(x)S_{\rho_{B}}(x)$. A particular set of combinations of addition and multiplication of two non-commutative variables that are relevant for our purposes is the squared commutator $(-i)^{2}\big{[}A(t),B(0)\big{]}^{2}$ and its higher-order generalizations. If $A(t)$ and $B(0)$ are asymptotically free, a prediction for the spectral statistics of this operator would help us to understand, in precise terms, the connection between previous proposals regarding the spectral statistics of out-of-time-order operators and the Lyapunovian operator rozenbaum2019universal  [^19] We have reviewed this proposal briefly in appendix E., and the emergence of asymptotic freeness in chaotic quantum many-body dynamics that we have advocated here. [^20] We also note here that, it is known in the free probability literature how to obtain the eigenvalue density of the commutator of two operators from the individual eigenvalue density nica1998commutators . One can follow such an approach to derive the free probability prediction for the convolved distribution of the commutator of, say, two spin-$1/2$ operators and check whether it matches the results from numerical simulations for interacting quantum many-body systems.
@@
  • It is crucial to determine the precise position of the notion of freeness (F-property) within the quantum ergodic hierarchy. We hypothesize that it lies between mixing and $K$-mixing, but further investigation is required to determine whether this is indeed true. Furthermore, exploring the relationship between $K$-mixing, the existence of quantum Lyapunov exponents, and stronger notions of freeness, such as second-order freeness collins2024second , is an important direction for future research.
@@ Acknowledgements.
We would like to thank Antonio Garcia-Garcia, Ping Gao, and Zhenbin Yang for valuable discussions on preliminary versions of this work. We also thank Pratik Nandy, Tanay Pathak, Rathindra Nath Das, and Leo Shaposhnik for useful discussions and comments on the draft. We also thank Pratik Nandy for carefully checking many numerical calculations in the draft. The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University, where this work was presented and discussed during the YITP workshop Quantum Gravity and Information in an Expanding Universe (YITP-W-24-19). This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\&& Future Planning (NRF-2021R1A2C1006791), by the Ministry of Education (NRF-2020R1I1A2054376) and the AI-based GIST Research Scientist Project grant funded by the GIST in 2025. This work was also supported by Creation of the Quantum Information Science R&\&&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). H. A. Camargo, and V. Jahnke were supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). H. Camargo, Y. Fu, V. Jahnke, and K. Pal should be recognized as co-first authors.
@@ Appendix A Cumulants from partitions
$\mathbb{E}(x_{1}x_{2}\cdots x_{n})=\sum_{\pi\in P(n)}c_{\pi}(x_{1}x_{2}\cdots x_{n}),~{}~{}\text{where}~{}~{}c_{\pi}(x_{1}x_{2}\cdots x_{n})=\prod_{b\in\pi}c_{|b|}\Big{(}x_{b(1)}x_{b(2)}\cdots x_{b(n)}\Big{)}~{}.$ (69)
Non-commutating random variables: Free cumulants. Now we consider the cumulants for non-commutating random variables ($A_{1},A_{2},\cdots A_{n}$), known as the free cumulants ($\kappa_{n}$). To relate these with $\phi(A^{n})$ we need to consider the partitions of the set $\{1,2,\cdots n\}$ which do not cross. While the total number of partitions of this set is given by the Bell numbers, the number of non-crossing partitions ($NC(n)$) is given by the Catalan numbers $C_{n}=\frac{1}{n+1}\Big{(}\begin{array}[]{c}2n\\ n\end{array}\Big{)}$, and the relation between $\phi(A^{n})$ and $\kappa_{n}$ of different order can be written as a sum over the non-crossing partitions in the following way,
@@
Here $b$ denotes the different elements of the blocks of the partitions, $b=(b(1),b(2),\cdots b(n))$ and $|b|$ indicates its length.
@@ Appendix D Average correlators and spectral form factor
To obtain the $2k$-point form factor from an OTOC average, one can consider the following average of a $2k$-point OTOC, and using the first moment of the Haar ensemble $12k-$ 1 times obtain Cotler:2017jue ,
@@
To summarize, these relations connecting the averages of the time-ordered correlators and OTOC with various spectral form factors indicate that the spectral statistics can not capture either of the spatial and temporal locality of the operators.
Operator averages for random matrix Hamiltonians. The relations between the averages of the correlation function and the SFFs discussed above are valid for any arbitrary Hamiltonian. One can also derive similar such relations when the Hamiltonian is drawn from some random matrix ensemble and the average is performed over the ensemble. For instance, for a random matrix Hamiltonian taken from the GUE, utilizing the invariance of the integral measure under unitary transformations, the following relation can be established between the ensemble-averaged 2-point correlation function and the 2-point form factor of the ensemble Cotler:2017jue
@@
If $\mathcal{U}$ is a non-identity Pauli operator, and $\mathcal{R}_{2}(t)\gg 1$, the above formula implies a relation between the GUE averaged 2-point function and 2-point form factor for the ensemble, which is very similar to the one in eq. (87),
@@
Computing the $2k$-point OTOC for the GUE one can similarly show that
$\big{<}\mathcal{U}_{1}(0)\mathcal{V}_{1}(t)\cdots\mathcal{U}_{k}(0)\mathcal{V}_{k}(t)\big{>}_{\text{GUE}}\simeq\big{<}\mathcal{U}_{1}\mathcal{V}_{1}\cdots\mathcal{U}_{k}\mathcal{V}_{k}\big{>}_{\text{GUE}}\frac{\mathcal{R}_{2k}(t)}{L^{2k}}~{}.$ (94)
This is non-zero only when $\mathcal{U}_{1}\mathcal{V}_{1}\cdots\mathcal{U}_{k}\mathcal{V}_{k}=I$, and in that case, this equation resembles the one in eq. (91). Performing a similar analysis one can show that
$\big{<}\mathcal{U}_{1}(0)\mathcal{U}_{2}(t)\mathcal{U}_{3}(2t)\mathcal{U}_{4}(t)\big{>}_{\text{GUE}}=\big{<}\mathcal{U}_{1}(0)\mathcal{U}_{2}(t)\mathcal{U}_{3}(0)\mathcal{U}_{4}(t)\big{>}_{\text{GUE}}\,$
@@
Appendix E Universal statistics of the out-of-time order operators
@@
Since the dependence of the OTOC on the state with respect to which it is computed is not universal, rozenbaum2019universal proposed to instead look at statistics of the so-called Lyapunovian operator (which is a Hermitian operator obtained by taking the logarithm of the squared commutator of two operators) to search for random matrix universality in these correlators. Below we first review this proposal more precisely. Consider the following correlators between two Hermitian operators, and define the operators $\mathbf{\Lambda}_{k}$,
@@
According to the proposal advocated in rozenbaum2019universal , one first construct the $N\times N$ matrices $\mathcal{C}^{k}_{nm}(t)=\mbox{$\langle E_{n}|$}\mathcal{C}^{k}(t)\mbox{$|E_{m}\rangle$}$ and $\mathcal{F}_{nm}(t)=\mbox{$\langle E_{n}|$}\mathcal{F}(t)\mbox{$|E_{m}\rangle$}$, where $|E_{m}\rangle$ denotes the energy eigenstates. Subsequently, these matrices are diagonalized by taking $t$ to be a fixed value, and the statistics of the spacing between the eigenvalues (as well as the logarithm of the eigenvalues) are to be obtained to look for random matrix universality in the correlators. One can then repeat this procedure for different values of time to compare the statistics as time evolves. This proposal was studied in detail in rozenbaum2019universal by taking the operators $A=x$, and $B=p_{x}$, and the Hamiltonian that of a stadium billiard

### v3

@@ 8.1 Future directions
  • In this work, we have examined the free convolution prediction for the sum of two free operators, focusing on spin-1/2 and spin-1 operators. While these results, specially, the spin-1/2 case, apply to a broad class of quantum mechanical systems, including spin chains and SYK-like models, it is crucial to extend the free convolution prediction to sums of other types of operators that naturally arise in quantum mechanics. In Sec. 4.1.1, we numerically determine the free additive convolution for the sum of two free spin-3/2, spin-2, spin-5/2, as well as large spin operators as a first step in this direction. However, many important cases remain unexplored, such as bosonic operators in, say, the Bose-Hubbard model, and the position $x$ and momentum $p$ operators in quantum billiard systems, and numerous others. A systematic classification of free convolution predictions for different classes of operators would be highly valuable for future studies on quantum chaos and asymptotic freeness in quantum mechanics.
@@
  • In this paper we have considered spectral properties of the sum of two non-commutative operators (say $A$ and $B$). However, one can similarly consider the spectral properties of the products of two such operators, as well as their different linear combinations. If these variables are traceless, i.e., $\varphi(A)=\varphi(B)=0$, then their free product is trivial since the trace of a combination of them vanishes from the definition freeness, $\varphi\big{(}(AB)^{k}\big{)}=\varphi\big{(}ABAB\cdots AB\big{)}=0$. For variables with non-zero trace, one can obtain the free multiplicative convolution of these variables by using an analogue of the $R$-transform, known as the $S$-transform, which has the property that for a pair of free variables, it is multiplicative voiculescu1987multiplication . This allows us to compute the free multiplicative convolution of the eigenvalue density of these two operators and obtain the eigenvalue density of the product by using the rule $S_{\rho_{A}\boxtimes\rho_{B}}(x)=S_{\rho_{A}}(x)S_{\rho_{B}}(x)$. A particular set of combinations of addition and multiplication of two non-commutative variables that are relevant for our purposes is the squared commutator $(-i)^{2}\big{[}A(t),B(0)\big{]}^{2}$ and its higher-order generalizations. If $A(t)$ and $B(0)$ are asymptotically free, a prediction for the spectral statistics of this operator would help us to understand, in precise terms, the connection between previous proposals regarding the spectral statistics of out-of-time-order operators and the Lyapunovian operator rozenbaum2019universal  [^27] We have reviewed this proposal briefly in appendix F., and the emergence of asymptotic freeness in chaotic quantum many-body dynamics that we have advocated here. [^28] We also note here that, it is known in the free probability literature how to obtain the eigenvalue density of the commutator of two operators from the individual eigenvalue density nica1998commutators . One can follow such an approach to derive the free probability prediction for the convolved distribution of the commutator of, say, two spin-$1/2$ operators and check whether it matches the results from numerical simulations for interacting quantum many-body systems.
@@
  • Exploring the emergence of stronger notions of freeness, such as second-order freeness collins2024second , under chaotic dynamics is also an interesting direction for future research.
@@ Acknowledgements.
We would like to thank Antonio Garcia-Garcia, Ping Gao, Zhenbin Yang, Neil Dowling, and Silvia Pappalardi for valuable discussions. We are also grateful to Pratik Nandy, Tanay Pathak, Rathindra Nath Das, Sergio E. Aguilar-Gutierrez, and Leo Shaposhnik for useful discussions and comments on the draft. In particular, we thank Pratik Nandy for carefully checking many of the numerical calculations. We also thank Heide Narnhofer for helpful correspondence. The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University, where this work was presented and discussed during the YITP workshop Quantum Gravity and Information in an Expanding Universe (YITP-W-24-19). This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\& Future Planning (NRF-2021R1A2C1006791), by the Ministry of Education (NRF-2020R1I1A2054376) and the AI-based GIST Research Scientist Project grant funded by the GIST in 2025. This work was also supported by Creation of the Quantum Information Science R&\&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). H. A. Camargo, and V. Jahnke were supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). H. Camargo, Y. Fu, V. Jahnke, and K. Pal should be recognized as co-first authors.
@@ Appendix A Cumulants from partitions
$\mathbb{E}(x_{1}x_{2}\cdots x_{n})=\sum_{\pi\in P(n)}c_{\pi}(x_{1}x_{2}\cdots x_{n})~,~~\text{where}~~~c_{\pi}(x_{1}x_{2}\cdots x_{n})=\prod_{b\in\pi}c_{|b|}\Big{(}x_{b(1)}x_{b(2)}\cdots x_{b(n)}\Big{)}~.$ (92)
Non-commutating random variables: Free cumulants. Now we consider the cumulants for non-commutating random variables ($A_{1},A_{2},\cdots A_{n}$), known as the free cumulants ($\kappa_{n}$). To relate these with $\varphi(A^{n})$, we need to consider the partitions of the set $\{1,2,\cdots n\}$ which do not cross. While the total number of partitions of this set is given by the Bell numbers, the number of non-crossing partitions ($NC(n)$) is given by the Catalan numbers $C_{n}=\frac{1}{n+1}\Big{(}\begin{array}[]{c}2n\\ n\end{array}\Big{)}$, and the relation between $\varphi(A^{n})$ and $\kappa_{n}$ of different order can be written as a sum over the non-crossing partitions in the following way,
@@
Here $b(I)$ denotes the different elements of the blocks of the partitions, $b=(b(1),b(2),\cdots b(n))$ and $|b|$ indicates its length.
@@ Appendix D Average correlators and spectral form factor
To obtain the $2k$-point form factor from an OTOC average, one can consider the following average of a $2k$-point OTOC, and using the first moment of the Haar ensemble 2k12k-1 times obtain Cotler:2017jue ,
@@
To summarize, these relations connecting the averages of the time-ordered correlators and OTOC with various spectral form factors indicate that the spectral statistics can not capture either of the spatial or temporal locality of the operators.
Average correlation function for random matrix Hamiltonians. The relations between the averages of the correlation function and the SFFs discussed above are valid for any arbitrary Hamiltonian. One can also derive similar such relations when the Hamiltonian is drawn from some random matrix ensemble and the average is performed over the ensemble. For instance, for a random matrix Hamiltonian taken from the GUE, utilizing the invariance of the integral measure under unitary transformations, the following relation can be established between the ensemble-averaged 2-point correlation function of two self-adjoint operators and the 2-point spectral form factor of the ensemble Cotler:2017jue
@@
Here $\mathcal{R}_{2}(t)$ is the spectral form factor defined as
$\mathcal{R}_{2}(t)=\int\text{d}H~\big{|}\text{Tr}\big{(}e^{-itH}\big{)}\big{|}^{2}~,$ (116)
and $\langle A\rangle=\text{Tr}(A)/L$. If $\mathcal{U}$ is a non-identity Pauli operator, and $\mathcal{R}_{2}(t)\gg 1$, the above formula implies a relation between the GUE averaged 2-point function and 2-point form factor for the ensemble, which is very similar to the one in eq. (110),
@@
Computing the $2k$-point OTOC for the GUE, one can similarly show that
$\big{<}\mathcal{U}_{1}(0)\mathcal{V}_{1}(t)\cdots\mathcal{U}_{k}(0)\mathcal{V}_{k}(t)\big{>}_{\text{GUE}}\simeq\big{<}\mathcal{U}_{1}(0)\mathcal{V}_{1}(0)\cdots\mathcal{U}_{k}(0)\mathcal{V}_{k}(0)\big{>}\frac{\mathcal{R}_{2k}(t)}{L^{2k}}~.$ (118)
This is non-zero only when $\mathcal{U}_{1}\mathcal{V}_{1}\cdots\mathcal{U}_{k}\mathcal{V}_{k}=I$, and in that case, this equation resembles the one in eq. (114). Performing a similar analysis, one can show that
$\big{<}\mathcal{U}_{1}(0)\mathcal{U}_{2}(t)\mathcal{U}_{3}(2t)\mathcal{U}_{4}(t)\big{>}_{\text{GUE}}\simeq\big{<}\mathcal{U}_{1}(0)\mathcal{U}_{2}(t)\mathcal{U}_{3}(0)\mathcal{U}_{4}(t)\big{>}_{\text{GUE}}\,$~,
@@
Finite temperature correlators.
So far, we have only considered the infinite temperature correlation functions. However, one can follow the strategy described above to compute the finite temperature correlation function as well and see whether they can be related to the finite temperature SFF. To begin with, consider the correlation function defined through the standard thermal expectation value, $\langle\mathcal{O}\rangle_{\beta}=Z^{-1}(\beta)\text{Tr}(e^{-\beta H}\mathcal{O})$, $Z(\beta)$ being the partition function. For a random matrix belonging to the GUE, using the unitary invariance of the ensemble, we have [^30] In order to avoid cluttering the notation, we have removed the subscript GUE in the correlation functions used previously.
$\begin{split}\langle AB(t)\rangle_{\beta}=\int\text{d}U~\text{d}H~\langle A_{U}e^{iHt}B_{U}e^{-iHt}\rangle_{\beta}=\frac{\langle AB\rangle}{(L^{2}-1)}\big{(}L~r_{2}(t)-1\big{)}+\frac{L\langle A\rangle\langle B(t)\rangle}{L^{2}-1}\big{(}L-r_{2}(t)\big{)}~.\end{split}$ (119)
Here we have defined,
$r_{2}(t)=\frac{1}{Z(\beta)}\int\text{d}H~\text{Tr}\big{(}e^{-itH}\big{)}\text{Tr}\big{(}e^{iH(\beta+it)}\big{)}~,~~Z(\beta)=\int\text{d}E\sum_{j}e^{-\beta E_{j}}~.$ (120)
On the other hand, using the Wightman inner product instead of the standard thermal inner product, we have the following relation
$\langle AB(t)\rangle_{\text{W}}=\langle e^{\beta H/2}Ae^{-\beta H/2}e^{iHt}Be^{-iHt}\rangle_{\beta}\\ =\frac{L}{Z(\beta)}\int\text{d}U~~\langle A_{U}B_{U}\Big{(}t-\frac{i\beta}{2}\Big{)}\rangle~.$ (121)
Following a procedure similar to the one in (28), we get the expression for the Wightman two-point correlation function at an inverse temperature $\beta$ to be
$\langle AB(t)\rangle_{\text{W}}=\frac{L}{Z(\beta)}\Bigg{[}\langle AB\rangle\bigg{(}\frac{L\mathcal{R}_{2}(t,\beta/2)-Z(\beta)}{L(L^{2}-1)}\bigg{)}+\langle A\rangle\langle B\rangle\bigg{(}\frac{LZ(\beta)-\mathcal{R}_{2}(t,\beta/2)}{(L^{2}-1)}\bigg{)}\Bigg{]}~,$ (122)
where we have defined analytically continued partition function $Z(t,\beta/2)=\text{Tr}\big{(}e^{-iH(t-i\frac{\beta}{2})}\big{)}$ and the associated SFF, $\mathcal{R}_{2}(t,\beta/2)=\langle Z(t,\beta/2)Z^{*}(t,\beta/2)\rangle_{\text{GUE}}$.
Appendix E Towards operator statistics at finite temperature
As we have discussed in section 8, the predictions for the operator statistics obtained from the free probability theory that we have used in the main text are applicable for a quantum system at infinite temperature. Here we provide a brief discussion on how one can proceed in obtaining the statistics of the sum of free operators for finite temperature states.
We begin by noticing that, in general, if two operators $A(0)$ and $B(t)$ are free with respect to an infinite-temperature state, we then expect them to also be approximately free with respect to finite-temperature states, as long as the temperature is sufficiently high. However, as the temperature is lowered, we approach the regime where expectation values are computed near the quantum system’s ground state. In this case, we do not generally expect freeness, except perhaps in special systems that remain highly chaotic even near the ground state.
The free probability prediction for the free convolution of two operators is obtained by computing the Cauchy transform $G(z)$, whose definition is appropriate for an infinite-temperature state (see eq. (36) and the discussion in footnote 17). One approach to obtaining the free convolution prediction for a finite temperature state with the map, taken e.g., the standard thermal inner product $\varphi_{\beta}(\mathcal{O})=\langle\mathcal{O}\rangle_{\beta}=\text{Tr}(\rho_{\beta}\mathcal{O})/\text{Tr}(\rho_{\beta})$, $\rho_{\beta}=e^{-\beta H}$ being the thermal density matrix, would be to use an appropriate generalization of the Cauchy transform which takes into account the finite temperature of the quantum system.
However, in principle, one can choose a different definition of the thermal inner product as well. In particular, choosing the Wightman inner product to define the thermal two-point function, we see that for a generic Hamiltonian $H$, it can be rewritten in terms of an infinite temperature state as
$\langle A(0)B(t)\rangle_{\text{W}}=\langle\rho_{\beta}^{1/2}A\rho_{\beta}^{-1/2}e^{iHt}Be^{-iHt}\rangle_{\beta}=\frac{L}{Z(\beta)}\langle A(0)B\Big{(}t-\frac{i\beta}{2}\Big{)}\rangle~.$ (123)
Therefore, this two-point function also captures thermal effects and is related to the spectral properties of $A(0)+\rho^{1/2}B(t)\rho^{1/2}$. By considering the composite operator $\rho^{1/2}B\rho^{1/2}$ (instead of just $B$), we can incorporate thermal effects into the operator $B$ while still using the (infinite temperature) map $\varphi(\cdot)=\text{Tr}(\cdot)/N$. This allows us to apply the formalism of the standard Cauchy transform and obtain the free convolution prediction, while still incorporating thermal effects. The price to pay is that the spectrum of $\rho^{1/2}B\rho^{1/2}$ differs from that of $B$ and needs to be determined.
To this end, one could proceed as follows. If the dynamics of a quantum system is chaotic and Hamiltonian ($H$) behaves like a random matrix, we expect $\rho^{1/2}=e^{-\beta H/2}$ to also approximately behave like a random matrix and to be free from the deterministic operator $B$. This allows us to use free multiplicative convolution, based on the S-transform, to find the spectrum of $\rho^{1/2}B\rho^{1/2}$. Knowing the spectra of $A(0)$ and $\rho^{1/2}B\rho^{1/2}$, we can compute the corresponding R-transforms and thus determine the spectral properties of $A(0)+\rho^{1/2}B\rho^{1/2}$. However, the details of this procedure are beyond the scope of the present paper, and we plan to report on finite-temperature generalizations of our results in the near future.
Appendix F Universal statistics of the out-of-time order operators
@@
Since the dependence of the OTOC on the state with respect to which it is computed is not universal, rozenbaum2019universal proposed to instead look at statistics of the so-called Lyapunovian operator (which is a Hermitian operator obtained by taking the logarithm of the squared commutator of two operators) to search for random matrix universality in these correlators. Below, we first review this proposal more precisely. Consider the following correlators between two Hermitian operators, and define the operators $\mathbf{\Lambda}_{k}$,
@@
According to the proposal advocated in rozenbaum2019universal , one first constructs the $N\times N$ matrices $\mathcal{C}^{k}_{nm}(t)=\mbox{$\langle E_{n}|$}\mathcal{C}^{k}(t)\mbox{$|E_{m}\rangle$}$ and $\mathcal{F}_{nm}(t)=\mbox{$\langle E_{n}|$}\mathcal{F}(t)\mbox{$|E_{m}\rangle$}$, where $|E_{m}\rangle$ denotes the energy eigenstates. Subsequently, these matrices are diagonalized by taking $t$ to be a fixed value, and the statistics of the spacing between the eigenvalues (as well as the logarithm of the eigenvalues) are obtained to look for random matrix universality in the correlators. One can then repeat this procedure for different values of time to compare the statistics as time evolves. This proposal was studied in detail in rozenbaum2019universal by taking the operators $A=x$, and $B=p_{x}$, and the Hamiltonian that of a stadium billiard
@@
so that (in units $\hbar=1$) $\mathcal{F}(t=0)=-\Big{(}x\big{[}H,x\big{]}\Big{)}^{2}+\text{H.c}$. At the initial times, the bulk level spacing statistics of the operators $\mathcal{F}(0)$ and $\Lambda(0)$ are found to be the same as that of the Hamiltonian itself, i.e., they are closely approximated by the GOE Wigner-Dyson distribution. [^31] There are subtleties from different parity sectors of the eigenstates and the so-called bouncing ball modes that one needs to take care of before obtaining the level spacing statistics. See rozenbaum2019universal for more details. On the other hand, the level spacing distribution of these operators is different when the time takes a non-zero value. For finite values of time and odd values of k=2n1k=2n-1, the bulk level spacing statistics of $\mathcal{C}^{k=2n-1}(t\neq 0)$ follow GUE Wigner-Dyson distribution, while for even values of $k=2n$, the bulk level spacing of statistics of every second eigenvalue of the matrix $\mathcal{C}^{k=2n}_{ij}(t\neq 0)$ matches with the level spacing distribution of a Gaussian ensemble with the Dyson index $\nu=3$.
Appendix G Example of subordination function notebook
In this section, we provide a Mathematica code example that implements the subordination function method speicher2019lecture , as reviewed in Section 4.1.1, to compute the free additive convolution of two random variables whose auxiliary functions (58) are identical $H_{a}(z)=H_{b}(z)$. This code is used to generate the free probability predictions shown in Figures 9 and 10.
Table 2: Mathematica code for free convolution based on the subordination function method
(* Solves fixed-point equation for $\omega(z)$ using 20,000 iterations *)
Clear[MapF]
MapF[z_] := Module[{HH, MAP, w0, wf},
   HH[x_] := Ha[x]; (* auxiliary function *)
   MAP[ww_] := z + HH[z + HH[ww]];
   w0 = 1. + 1.1 I;
   wf = Nest[MAP, w0, 20000];
   wf
(* Inverse Stieltjes transform *)